Required information A 1.5-in. steel plate and a 0.75-in. gray cast iron plate are compressed with one bolt and nut. The boltPLEASE PLEASE BE SURE THE ANSWER IS RIGHT, THANKS.
Answer #1

Date: Page no: sob From table A-32 Basic dimensions of American Standa&d plain washers, FOC Fastenes size of IN The washerDate: Page no: 2 The thread length is, L1 - 20 + 1/4 in = 2(0.5) +0-25 Lt = 1.25 in The length of unthreaded portions L-LT =